OpenStudy (haleyelizabeth2017):

A hyperbola has vertices (+-5, 0) and one focus (6, 0). What is the standard form equation of the hyperbola? I just would like some help starting....please no direct answers :)

OpenStudy (anonymous):

Ok! So what is standard form of a hyperbola?

OpenStudy (haleyelizabeth2017):

For this one, wouldn't it be \[(y-k)^2=4p(x-h)\] because it is opens up horizontally?

OpenStudy (anonymous):

Hm. Are you sure that's the standard equation?

OpenStudy (haleyelizabeth2017):

No...actually I am not... :(

OpenStudy (anonymous):

I think you are mistaken that is for a parabola

OpenStudy (haleyelizabeth2017):

oh yeah! lol one second

OpenStudy (anonymous):

We are looking for a hyperbola :)

OpenStudy (haleyelizabeth2017):

Okay one second lol I can't recall that formula off the top of my head right this second :(

OpenStudy (anonymous):

Do you need me to write it for you?

OpenStudy (haleyelizabeth2017):

I got it lol I just had a duh moment

OpenStudy (haleyelizabeth2017):

\[\frac{ (x-h)^2 }{ a^2 }+ \frac {(y-k)^2}{b^2}=1\] right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So your vertex is (h, k + a)

OpenStudy (haleyelizabeth2017):

Okay

OpenStudy (anonymous):

Jk

OpenStudy (anonymous):

It's (h + or - a , k)

OpenStudy (haleyelizabeth2017):

Okie dokie

OpenStudy (anonymous):

Do you know what a is now?

OpenStudy (haleyelizabeth2017):

5?

OpenStudy (anonymous):

Yup! Do you know what h and k are yet?

OpenStudy (haleyelizabeth2017):

(0, 0)? So the center is the origin....

OpenStudy (anonymous):

Yes! Now we just need to know what b is.

OpenStudy (haleyelizabeth2017):

b is -5, correct?

OpenStudy (anonymous):

Nope!

OpenStudy (haleyelizabeth2017):

:( I'm not sure then.....

OpenStudy (anonymous):

Dont worry! We can find b by a^2 + b^2 = c^2. c is part of the focus, (h + or - c, k)

OpenStudy (haleyelizabeth2017):

Okay....

OpenStudy (anonymous):

So lets fill in the values to find b. a^2 + b^2 = c^2 5^2 + b^2 = 6^2

OpenStudy (haleyelizabeth2017):

ah hah!

OpenStudy (haleyelizabeth2017):

so b^2 is going to be 11....so \[\frac{x^2}{25} + \frac{y^2}{11}=1\] right? I'm pretty sure that is it.....

OpenStudy (anonymous):

Yeah checks out

OpenStudy (haleyelizabeth2017):

Thank you!

OpenStudy (anonymous):

np

OpenStudy (haleyelizabeth2017):

@swagmaster47 that doesn't work because it then forms an ellipse....

OpenStudy (haleyelizabeth2017):

I'm so confused.....

OpenStudy (haleyelizabeth2017):

I figured out the problem

OpenStudy (anonymous):

You put a plus instead of a subtraction symbol between the two

OpenStudy (haleyelizabeth2017):

it's supposed to be plus.....but we solved for b incorrectly.

OpenStudy (haleyelizabeth2017):

it was supposed to be a^2=b^2+c^2....

OpenStudy (anonymous):

No we did b right

OpenStudy (haleyelizabeth2017):

so then 11 would be -11......I think it works either way...

OpenStudy (anonymous):

Cant be -11

OpenStudy (anonymous):

You squared b

OpenStudy (haleyelizabeth2017):

squared b? when did we square it?

OpenStudy (haleyelizabeth2017):

because it just cancels out when you square a square root.......so -11 would just be plugged in for b^2.....

OpenStudy (haleyelizabeth2017):

I am so lost :(

OpenStudy (haleyelizabeth2017):

it can't be a^2+b^2=c^2 because that is the pythagorean theorem..

OpenStudy (anonymous):

Technically b is sqrt(11)

OpenStudy (haleyelizabeth2017):

that is true...

OpenStudy (anonymous):

"The vertices are a units from the center, and the foci are c units from the center, with c^2 = a^2 + b^2. Notice that in the equations for a hyperbola, unlike an ellipse, a^2 is not necessarily larger than b^2"

OpenStudy (haleyelizabeth2017):

I'm so confused :(

OpenStudy (anonymous):

https://www.desmos.com/calculator/wd0kkajkqu

OpenStudy (anonymous):

You got it, but you put a plus between your fractions instead of a minus